OpenStudy (anonymous):

bella wants to write two equations to model the streets on this map. She can use y=x-4 to describe Platte Way. Find the one absolute value equation to describe Martin Rd and Smith St.

OpenStudy (anonymous):

is there supposed to be a diagram?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):



OpenStudy (anonymous):

is smith street parallel to platte way and is marcus road perpendicular to smith street?

OpenStudy (anonymous):

nvm i think i can see tell the points from the drawing. hopefuly i read them right. smith st. cuts through (-1,-1) and (-2,-2) and the origin. the gradient is thus 1 the equation is thus y = x marcus rd cuts through (-5,0) and (-3,-3) so the gradient is -1.5 y-0 = -1.5(x+5) y = -1.5x -7.5